alones (alones):

HEPL _ANYBODY!! I need to check if i got this right!! 2x + y = 6

TheSmartOne (thesmartone):

??

TheSmartOne (thesmartone):

The question is...?

alones (alones):

on top

pooja195 (pooja195):

what is the question?

OpenStudy (alexandervonhumboldt2):

do you need to find slope or what?

alones (alones):

no i have to do it in a form A GENERAL FORM Ax + By + C = 0.

alones (alones):

2x - y - 6 = 0 this what i've got

OpenStudy (alexandervonhumboldt2):

just subtarct 6 from both sides: 2x+y-6=6-6 ???

alones (alones):

so it's wong?

alones (alones):

ugh geez the prevois was 2x + y - 6 = 0

OpenStudy (alexandervonhumboldt2):

i got 2x+y-6=0

alones (alones):

so ma prevois was right

OpenStudy (alexandervonhumboldt2):

why? 2x+y-6=0 is correct

alones (alones):

whaa??

alones (alones):

so i've mrs.humbolt x-x and y -y so the first one which i got the prevoius one which the one you said ewas right

OpenStudy (alexandervonhumboldt2):

omg

alones (alones):

what ??

OpenStudy (alexandervonhumboldt2):

you are making me laught to death

alones (alones):

so now i got 2x + y - 6 = 0 HERE

OpenStudy (alexandervonhumboldt2):

yes

alones (alones):

REALLLY IT IS!!1!!?1

TheSmartOne (thesmartone):

@AloneS = @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

no?

OpenStudy (alexandervonhumboldt2):

are you joking? ip check using all methods? we are not alts

TheSmartOne (thesmartone):

you talk the same ip check shows you both

OpenStudy (alexandervonhumboldt2):

thesmatone, Alones is not my alt

OpenStudy (alexandervonhumboldt2):

we are not talking same. i gonna have amod to ip check

alones (alones):

non i leave in america and he lives somewhere else

OpenStudy (alexandervonhumboldt2):

stop it

alones (alones):

oh GEEZE GUYS i just asked a question thx Alot @AlexandervonHumboldt2 you've he43lPeddfcdxd alott i appreciate it morede thst some one 1else1`

OpenStudy (alexandervonhumboldt2):

tso is thinking we are alts because you have my pic

OpenStudy (alexandervonhumboldt2):

alones is just my friend

alones (alones):

YEAH THAT'S

alones (alones):

all we need to kn`

alones (alones):

so the answer is 2x + y - 6 = 0